3. A series combination of two uncharged capacitors are connected to a 15 V battery, 300 μJ of energy is drawn from the battery. If one of the capacitors has a capacitance of 5.0 pF, what is the capacitance of the other?


1
Expert's answer
2022-03-13T18:48:51-0400

The total capacitance "C" of the series can be found from the formula:


"E = \dfrac{CU^2}{2}"

where "E = 300\times 10^{-6}J" is the energy of the system, "U = 15V" is the voltage. Thus, obtain:


"C = \dfrac{2E}{U^2}"

On the other hand, the capacitance of the two capacitors in cseries is given as follows:


"\dfrac{1}{C} = \dfrac{1}{C_1} + \dfrac{1}{C_2}"

where "C_1 = 5\times 10^{-12}F, C_2" are individual capacitances. Expressing "C_2" and substituting the formula for "C", obtain:


"C_2 = \dfrac{CC_1}{C_1-C} = \dfrac{C_1}{C_1/C-1} = \dfrac{C_1}{C_1\dfrac{U^2}{2E}-1}\\
C_2 = \dfrac{5\times 10^{-12}F}{5\times 10^{-12}F\cdot \dfrac{(15V)^2}{2\cdot 300\times 10^{-6}J}-1} \approx -5\times 10^{-12}F"

The ansver is negative, thus, the given numbers are wrong.


Answer. "C_2 = \dfrac{C_1}{C_1\dfrac{U^2}{2E}-1}"